You manage the shipping of item #A452 from your supplier. The shipments for this item are delivered to the nearest port, and you have to transport them to the distribution center. You have two options, truck load (TL) and less than truck load (LTL).

Your company needs 5000 units of #A452 per year and it is purchased at a price of $36 per unit. Each order costs $97. Your company uses a holding charge of 0.15, a cycle service level of 90% and 365 days per year for planning purposes.

The truck load (TL) option costs $1200 per load. It takes 3 days to deliver and each truck can carry up to 1000 items.

The less than truck (LTL) load option costs $2.9 per item. It takes 5 days to deliver.

You have three options,

OPTION A: to use the TL option and deliver full truck loads,

OPTION B: to use the TL option and deliver the economic order quantity

OPTION C: to use LTL option and deliver the economic order quantity.

For all following parts, consider transportation costs as an item cost unless otherwise specified. i.e., the total variable cost of #A452 in the Full truckload is $1200/1000items + $36/item = $37.2/item.

(Option B) Truckload with EOQ

If you order the EOQ calculated earlier (rounded to a whole number), what is the sum of average pipeline, cycle inventory, ordering, and purchase costs for option B? Consider transportation costs a part of the item cost. Please provide your answers rounded to the closest dollar.
Answer #1

Solution :

Given:

Annual demand D=5000 units

Order cost S=97$

Holding cost i=0.15

Cost of item c =$36

Total holding cost H = 0.15*36=$5.4

Daily Demand = 5000/365=13.70

OPTION A :

Q=1000 units

D= 5000 units

Lead Time L=3 days

Service level = 90 %

Given Annua)见mand D-5000 ums Ordu Cost S:$91 holding rot 0.15 cost ot item C= 981.2 Holdiną Cast: Daily Demands 50OP : lyC :l13835 2 x 91 ) + ( 411 x 5, 58) 2 1163 01 1163.43 $ 2 326.5 2 321 3c5 ell inventory 2 2

1)Total cost = ordering cost
Answer #2
Ordering cost